OpenStudy (anonymous):

can you help me with this for my exercises: I need to to show that: g(x)= (e(-3x+2sinx)/6) ---------------- (3+2cosx) has derivative g'(x)= - (4(sin^2)x -12sin x +5) (e(-3x+2sinx)/6) ----------------------------------- 6(3+2cosx)^2

OpenStudy (anonymous):

this is g ? \[g(x) = \frac{e^{\frac{1}{6}(-3x +2sinx)}}{3+2cosx}\]

OpenStudy (anonymous):

let let f(x) be a function of x if we have a function of a function eg g(f(x)) when we differentiate we get g'(f(x)).f'(x)

OpenStudy (anonymous):

if we have two functions u(x) and v(x) multiplied together: u(x)v(x) when we differentiate we get u'(x)v(x) + v'(x)u(x)

OpenStudy (anonymous):

lets take a look at your function g(x)

OpenStudy (anonymous):

we could split it up: \[g(x) = \frac{1}{3+2cosx} \times e^{\frac{1}{6}(3x + 2sinx)} = u(x) \times v(x)\] \[u(x) = \frac{1}{3+2cosx} \] \[v(x) =e^{\frac{1}{6}(3x + 2sinx)}\] so to differentiate we need: \[u'(x)v(x) +v'(x)u(x)\] \[( \frac{1}{3+2cosx} \times \frac{d}{dx} (e^{\frac{1}{6}(3x + 2sinx)}) )+ (e^{\frac{1}{6}(3x + 2sinx)} \times \frac{d}{dx} (\frac{1}{3+2cosx}))\]

OpenStudy (anonymous):

ok so quotient rule says if h(x) = f(x)/g(x) then h'(x) = (f'(x)g(x) - g'(x)f(x))/(g(x))^2

OpenStudy (anonymous):

that is how we differentiate u(x) to differentiate v(x) we need to use the chain rule, which i started with :)

OpenStudy (anonymous):

Could you help do it with the quotient rule?

OpenStudy (anonymous):

Ok thanks.

OpenStudy (anonymous):

Appreciate your help